Where do i start to set this up?
Suppose that 79,000 is invested at 5 and 1/2% interest, compounded quarterly .
(a) find the function for the amount to which the investment grows after t years.
(b) find the amount of money in the account at t=0,3,7,10 years.
(c)when will the amount of money reach $300,000?

Answer by stanbon(75887)

Suppose that 79,000 is invested at 5 and 1/2% interest, compounded quarterly .
(a) find the function for the amount to which the investment grows after t years.
---
A(t) = P(1+(r/n))^(nt)
A(t) = 79000(1+(0.055/4))^(4t)
-------------------------------------
(b) find the amount of money in the account at t=0,3,7,10 years.
Substitute each of those numbers for "t" to find A(t)
--------------------------------------
(c)when will the amount of money reach $300,000?
Solve: 300,000 = 79000(1+(0.055/4))^(4t)
3.7975 = (1.01375)^(4t)
---
Take the log of both sides to get:
4t*log(1.01375) = log(3.7975)
4t = 97.71
time = 24.4 years
======================
Cheers,
Stan H.